# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

### v1

@@ 4.2 Expected Values and Variances Comparison of Premium Estimators
The expected value and variance of premium estimator in offset and ratio approach are obtained as follows:
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.4)
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.5)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]\right)^{2};$ (4.6)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]\right)^{2}.$ (4.7)
@@
The key to the proof is to note that, for a random variable $Z$, the relationship $\log(Z)\sim\mathcal{N}(\mu,\sigma^{2})$ holds. The expected value and variance of $Z$ are then given by:
  • $\mathrm{E}\!\left[Z\right]=\exp\{\mu+\frac{1}{2}\sigma^{2}\}$;
  • $\mathrm{Var}\!\left[Z\right]=\left(\exp\left\{\sigma^{2}\right\}-1\right)\left(\mathrm{E}\!\left[Z\right]\right)^{2}$.
Additionally, the asymptotic properties of the maximum likelihood estimator (MLE) allow us to derive the distributions of the parameter vector estimators as follows:
$\bm{\widehat{\beta}^{O}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \bm{\widehat{\beta}^{R}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\right).$
So, using these asymptotic distributions and Equations (4.2) and (4.2), we obtain the following relationships:
$\log\left(\bm{\widehat{\zeta}^{O}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \log\left(\bm{\widehat{\zeta}^{R}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right).$
This demonstrates how the expected values and variances of the premium estimators can be derived using the distributions of the parameter vector estimators. ∎
@@
The expected value of the premium estimator in offset approach is smaller than in the ratio approach. Additionally, the variance of the premium estimator in the offset is also smaller than in the ratio approach. Formally, for contract $i$, we obtain:
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right];$ (4.8)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right].$ (4.9)
@@
Referring to the expressions for the expected values and variances of premium estimators given in Proposition 1, we observe that the key difference between these metrics arises from the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regressions: $\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}$.
This proof, therefore, focuses on comparing these covariance matrices between the two approaches. As noted in Section 2.4.1, both covariance matrices are well defined and positive definite:
$\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\in\mathcal{P}_{q+1}.$
Let us define the matrix $M$ as the difference between these two covariance matrices:
$M=\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}-\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}.$
The goal is to show that $M$ is a positive definite matrix. To do so, we can utilize the positive definite matrix property described in Section 2.8. Therefore, proving that $D^{O}-D^{R}$ is a positive definite matrix is sufficient to show that $M$ is also one.
By examining Figure 1, we observe that $D^{O}-D^{R}$ is indeed a positive definite matrix. This can be formally expressed as:
$\forall v=(v_{1},\ldots,v_{n})^{t}\in\mathbf{R}^{n}:v^{t}\left(D^{O}-D^{R}\right)v=\sum_{i=1}^{n}\left(t^{2-p}_{i}-t_{i}\right)\zeta^{2-p}_{i}v^{2}_{i}>0.$
Additionally, using equation (2.9), we note that the following matrix is positive definite because the matrix $\bm{X}$ has full rank $q+1$:
$\bm{X}^{t}\left(\bm{D}^{O}-\bm{D}^{R}\right)\bm{X}=\bm{X}^{t}\bm{D}^{O}\bm{X}-\bm{X}^{t}\bm{D}^{R}\bm{X}.$
Thus $M$ being positive definite implies that the expected value and variance of the premium estimator in the offset approach are smaller than in the ratio approach, as stated in Proposition 2. ∎
From equations (4.4) and (4.5), it is evident that there is a positive bias in the premium estimator for both regression approaches, as $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]-\zeta_{i}>0$ and $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]-\zeta_{i}>0$. However, using Proposition 2, we observe that the bias of the premium estimator in the ratio approach is consistently larger than that in the offset approach. Therefore, based on the bias and variance criterion, the premium estimator from the offset approach is superior to that of the ratio approach.
It is important to note that, in insurance pricing, the primary focus is on the bias of the premium estimator. Indeed, the goal for an insurer is to develop a pricing model that significantly minimizes premium estimator bias. Auto-calibration is generally used for this purpose. For further details on premium estimator bias reduction, recent works by Denuit et al. (2021), Lindholm et al. (2023), Denuit et al. (2024) and Denuit and Trufin (2024) can be consulted. However, this paper prioritizes bias reduction at the portfolio level rather than at the individual premium level.
4.3 Random Gaps Comparison
Another comparison between the two approaches can be made based on what we will refer to as the random gaps. More formally, this refers to the following values for the offset and ratio approaches, respectively:
$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right).$
These expressions represent the difference between two random variables: one corresponding to the loss cost of insured individual $i$, and the other to the premium estimator for the same individual. We refer to the expectation of the random gap as the "portfolio bias". According to Denuit et al. (2024), financial equilibrium at the portfolio level corresponds to a zero portfolio bias. In other words, the expectation of the random gap should be close to zero to ensure financial equilibrium at the portfolio level.
However, due to the positive bias in the premium estimators for both regression approaches, financial equilibrium is not achieved in either the offset or ratio approach at the portfolio level. Using Proposition 2, we note that the portfolio bias in the ratio approach is smaller than in the offset approach. Formally, we establish the following relationship:
$\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\leq\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right].$ (4.10)
This relationship implies that the premium estimator from the ratio approach is more effective than that from the offset approach in achieving financial equilibrium across the entire portfolio. Our objective is to provide a probabilistic framework for comparing the full distributions of these random gaps using stochastic dominance. Specifically, first-order stochastic dominance allows us to demonstrate that the ratio approach results in smaller gaps with higher probability than the offset approach.
Before presenting our results, we first formally define stochastic dominance between two random variables, based on the work of Shaked and Shanthikumar (2007).
Definition 4.3.1.
Consider two random variables, $R_{1}$ and $R_{2}$. We say $R_{2}$ dominates $R_{1}$, denoted by $R_{1}\leq_{st}R_{2}$, if the following condition is satisfied:
$\mathrm{Pr}\!\left(R_{1}>x\right)\leq\mathrm{Pr}\!\left(R_{2}>x\right),\forall x\in\mathbf{R}.$
Additionally, we provide a necessary and sufficient condition for ordering two random variables, as outlined in Shaked and Shanthikumar (2007).
Property 4.3.1.
For two random variables, R1 and R2, the following relationship holds:
$R_{1}\leq_{st}R_{2}\Longleftrightarrow\mathrm{E}\!\left[f(R_{1})\right]\leq\mathrm{E}\!\left[f(R_{2})\right],$
where $f(\cdot)$ is an increasing function for which the expectations exist.
The two previous results enable us to state the following property:
Proposition 3.
The difference between the aggregate loss costs and the sum of the premium estimators is more likely to take a significant value in the offset regression approach than in the ratio regression approach. Formally, for contract $i$, we establish the following stochastic dominance:
$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\leq_{st}\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right).$ (4.11)
Proof.
The key to this proof is using the Taylor series approximation as discussed in Van der Vaart (2000). Specifically, the idea is to approximate the expected value of a function of a random variable by applying the function to the expected value of the random variable itself.
Consider a function $f(\cdot)$. By applying the first-order Taylor approximation to the function $f(\cdot)$ at the sums of the difference between the aggregate loss cost and the sum of the premium estimators, $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)$ and $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)$, we obtain the following results for each contract $i$:
$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right]\right);$ (4.12)
$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\right).$ (4.13)
Using Equation (4.10) and considering $f(.)$ as an increasing function, we conclude that:
$\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]\leq\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right].$
This proof shows that the gap between the aggregate loss costs and the sum of the premium estimators is more likely to be larger in the offset regression approach than in the ratio regression approach.
4.3.1 Remarks
In general, for each contract $i,i=1,\ldots,n$, when the premium, $\zeta_{i}$, is linked to the score $\mathbf{X}^{T}_{i}\beta$ as $\zeta_{i}=g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right)$, $\widehat{\beta}$ serves as the estimator of $\beta$ in both the offset and ratio approaches. However, except in the case of the identity link function, $s>0\mapsto g(s)=s$, the premium estimator $\zeta_{i}$ has non-zero asymptotic bias:
$\mathrm{E}\!\left[g^{-1}\left(\mathbf{X}^{T}_{i}\widehat{\beta}\right)\right]\neq g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right).$
Consequently, financial equilibrium based on the expected random gap cannot be achieved unless the identity link is used in premium modelling. For this reason, in GLM, insurers consider the balance property to ensure an alternative financial equilibrium based on the alignment of observed loss costs with estimated premiums (see Nelder and Wedderburn (1972) for a discussion on the relationship between the sum of the predicted values and the sum of the observed values in GLM theory). As discussed in Section 2, this financial equilibrium ensures that the sum of the estimated premiums equals the aggregate observed loss costs across the entire portfolio, as well as at each level of risk factors used in premium modelling.
It is important to note that even if the balance property is not verified in either approach, Proposition 3 indicates that it is more likely for the random gaps to take on larger values in the offset approach than in the ratio approach. However, this alone is insufficient to determine which approach—offset or ratio—is preferable for achieving financial equilibrium under the balance property. Therefore, we need to examine the difference between aggregate observed loss costs and the sum of estimated premiums for both approaches.
4.4 Observed Gaps Comparison
We refer to the observed gaps in each regression approach as defined below:
$\sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right),$ (4.14)

### v2

@@ 4.1 Parameter Vector Estimators Comparison
According to White (1982), if the Kullback–Leibler divergence between the true density $g$ and the density $f$ (assumed under either the offset or the ratio approach) exists, then the corresponding estimator of the true parameter vector is consistent under both approaches. More formally, if
$KL(\bm{\beta})<+\infty\quad\text{for all }\bm{\beta},$
then
$\widehat{\bm{\beta}}^{O}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}},\quad\widehat{\bm{\beta}}^{R}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}}\quad\text{as }n\to+\infty,$
where $\bm{\widehat{\beta}}^{O}$ and $\bm{\widehat{\beta}}^{R}$ denote the estimators of $\bm{\beta}^{\text{True}}$ obtained from the offset and ratio approaches, respectively.
@@
The function $KL(\cdot)$ is always nonnegative and finite whenever the probability measures induced by $g(\cdot\mid\bm{\beta}^{\text{True}},t)$ and $f(\cdot\mid\bm{\beta},t)$ are mutually absolutely continuous (see Gray (2011)). Since both $g(\cdot\mid\bm{\beta}^{\text{True}},t)$ and $f(\cdot\mid\bm{\beta},t)$ are densities, this condition is equivalent to requiring that their associated probability measures be mutually absolutely continuous, which follows from the Radon–Nikodym theorem.
Suppose that the minimum of $KL(\bm{\beta})$ exists and denote it by $\bm{\beta}^{*}$. According to White (1982), the estimators obtained under the working density $f$ converge in probability to this minimizer $\bm{\beta}^{*}$. Therefore, it suffices to show that this minimizer coincides with the true parameter vector, that is,
$\bm{\beta}^{*}=\bm{\beta}^{\text{True}}.$
We start from the following equivalence:
$\arg\min_{\bm{\beta}}KL(\bm{\beta})=\arg\max_{\bm{\beta}}\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$
For notational convenience, we write
$\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz\;=\;\int_{0}^{+\infty}\log(f)\,g\,dz.$
Using the expression of the Tweedie density, we obtain
$\int_{0}^{+\infty}\log(f)\,g\,dz$ $=\int_{0}^{+\infty}\Bigg[\frac{w}{\phi}\Bigg(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\Bigg)+a(z,w,\phi)\Bigg]g(z\mid\bm{\beta}^{\text{True}},t)\,dz$
$\propto\int_{0}^{+\infty}\Bigg(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\Bigg)g(z\mid\bm{\beta}^{\text{True}},t)\,dz,$
where the proportionality follows from the fact that $a(z,w,\phi)$ does not depend on $\bm{\beta}$.
By linearity of the integral, this expression becomes
$\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\int_{0}^{+\infty}z\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\int_{0}^{+\infty}g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$
Since
$\int_{0}^{+\infty}z\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz=\zeta^{\text{True}}=\exp\!\left\{\bm{x}^{\top}\bm{\beta}^{\text{True}}\right\},\qquad\int_{0}^{+\infty}g(z\mid\bm{\beta}^{\text{True}},t)\,dz=1,$
the objective function can be written as
$K(\bm{\beta})=\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}+\bm{x}^{\top}\bm{\beta}^{\text{True}}\}}{1-p}-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}.$
The gradient of $K(\bm{\beta})$ with respect to $\bm{\beta}$ is
$\nabla_{\bm{\beta}}K(\bm{\beta})=\Big(\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}+\bm{x}^{\top}\bm{\beta}^{\text{True}}\}-\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}\Big)\bm{x}.$
At the minimizer $\bm{\beta}^{*}$, the first-order condition yields
$\nabla_{\bm{\beta}}K(\bm{\beta}^{*})=\bm{0}\quad\Longleftrightarrow\quad\bm{\beta}^{*}=\bm{\beta}^{\text{True}},$
which completes the proof. ∎
4.2 Asymptotic Variance Approximation
Another important result established by White (1982) is that, under both the offset and ratio approaches, the estimators of the true parameter vector $\bm{\beta}^{\text{True}}$ are asymptotically Gaussian with a common mean equal to $\bm{\beta}^{\text{True}}$. Once consistency is established, differences between the two approaches therefore arise solely through their asymptotic dispersion rather than through their limiting value.
From a practical standpoint, differences in asymptotic dispersion translate into differences in the precision and stability of the estimated covariate effects. A smaller asymptotic covariance matrix implies tighter confidence intervals for the regression coefficients, which facilitates the comparison, selection, and validation of pricing variables, particularly when some effects are modest but economically relevant. At the same time, reduced dispersion limits the sensitivity of the estimated parameters to random portfolio fluctuations.
In this context, the comparison of asymptotic covariance matrices provides a natural and meaningful criterion for assessing the relative efficiency of the two estimation strategies. Using the asymptotic expansion derived in Appendix Appendix 1: Calculation of Asymptotic Covariance Matrices, we may therefore state the following proposition.
@@ 4.2 Asymptotic Variance Approximation
Consider the leading Gaussian terms of the Edgeworth expansion of the asymptotic covariance matrices of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$ associated with the offset and ratio approaches. Assume that these leading terms are given by
$\Sigma^{O}$ $=\phi(\bm{X}^{\top}\bm{D}^{O}\bm{X})^{-1},$ (4.1)
$\Sigma^{R}$ $=\phi(\bm{X}^{\top}\bm{D}^{R}\bm{X})^{-1}.$ (4.2)
Under the assumptions of Proposition 1 and for risk exposure satisfying $t\leq 1$, these leading Gaussian terms satisfy
$\Sigma^{R}-\Sigma^{O}\succ 0.$
@@
First, consider $n$ observations of the loss cost per unit of exposure $Z_{1},\ldots,Z_{n}$, with $Z_{i}\sim\operatorname{Tweedie}(\zeta_{i},w_{i},\phi,p)$, $w_{i}\in\{t_{i},t_{i}^{2-p}\}$, and $\zeta_{i}=\exp\{\bm{X}_{i}^{\top}\bm{\beta}\}$. According to White (1982), the asymptotic covariance matrix of the estimator is given by $\Sigma=A^{-1}BA^{-1}$, where $A=-I_{n}(\bm{\beta})$ and $B=\mathbb{E}[\nabla(\bm{\beta})\nabla(\bm{\beta})^{\top}]$, with
$\nabla(\bm{\beta})=\frac{1}{\phi}\Big(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,0},\ldots,\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,q}\Big)^{\top}.$
As shown in Appendix Appendix 1: Calculation of Asymptotic Covariance Matrices, $B=\phi^{-1}\bm{X}^{\top}\bm{D}\bm{X}=-A$, yielding
$\Sigma=A^{-1}BA^{-1}=-A^{-1}=\phi(\bm{X}^{\top}\bm{D}\bm{X})^{-1}.$
We now compare the two covariance matrices. As noted in Section 2.4.1, $\Sigma^{O},\Sigma^{R}\in\mathcal{P}_{q+1}$. Define $M=\Sigma^{R}-\Sigma^{O}$. To prove that $M$ is positive definite, it suffices, by the results of Equation (2.8), to show that $\bm{D}^{O}-\bm{D}^{R}$ is positive definite. From Figure 1, we have
$\forall v\in\mathbb{R}^{n},\quad v^{\top}(\bm{D}^{O}-\bm{D}^{R})v=\sum_{i=1}^{n}(t_{i}^{2-p}-t_{i})\zeta_{i}^{2-p}v_{i}^{2}>0.$
Moreover, since $\bm{X}$ has full rank $q+1$, equation (2.9) implies
$\bm{X}^{\top}(\bm{D}^{O}-\bm{D}^{R})\bm{X}=\bm{X}^{\top}\bm{D}^{O}\bm{X}-\bm{X}^{\top}\bm{D}^{R}\bm{X}\succ 0,$
It then follows from Equation (2.8) that $M=\Sigma^{R}-\Sigma^{O}$ is positive definite.
Taken together, the previous propositions show that each approach can be implemented independently and yields a consistent (i.e., convergent) estimator of the true parameter vector $\bm{\beta}^{\text{True}}$. However, based solely on the result stated in Proposition 2, one might be led to conclude that the estimator obtained under the offset approach is asymptotically more efficient than the one obtained under the ratio approach.
It is important to emphasize that Proposition 2 is formulated exclusively at the level of the leading Gaussian term of the Edgeworth expansion and therefore concerns only first-order asymptotic quantities. More specifically, the matrices $\Sigma^{O}$ and $\Sigma^{R}$ represent the leading $O(1/n)$ terms in the covariance expansions of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$. Consequently, the difference $\Sigma^{R}-\Sigma^{O}$ identified in Proposition 2 is itself a first-order quantity of order $O(1/n)$. The apparent superiority of the offset approach therefore relies on a comparison between leading terms only and is meaningful insofar as the full asymptotic covariance matrices of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$ are sufficiently well approximated by their leading Gaussian components $\Sigma^{O}$ and $\Sigma^{R}$, as defined in equations (4.1) and (4.2). Since higher-order contributions in the Edgeworth expansion are neglected, the resulting approximation error is of smaller order asymptotically, but may nevertheless be of the same order of magnitude as the first-order difference $\Sigma^{R}-\Sigma^{O}$ in finite samples. As a result, it is likely inappropriate to draw definitive conclusions regarding the relative performance of the two approaches based solely on Proposition 2. A natural extension would therefore consist in deriving and comparing higher-order terms in the Edgeworth expansion. While such an analysis would be of clear theoretical interest, it lies beyond the scope of the present paper.
4.2.1 Simulations Study
Instead, we focus on assessing the accuracy of the asymptotic approximation and on comparing the two methods through a simulation study. The simulation study is designed solely to assess the empirical relevance of Proposition 2, which provides a first-order asymptotic comparison between the offset and ratio approaches based on the leading Gaussian term of the Edgeworth expansion of the estimators’ covariance matrices. To this end, independent samples are generated under a data-generating process that is fully consistent with the modeling assumptions underlying the theoretical result. For each scenario, defined by a combination of sample size $n$ and Tweedie power parameter $p$, repeated Monte Carlo replications are performed, and both the offset and ratio estimators are computed using the same Tweedie specification as that used for data generation.
For each approach, the resulting collection of parameter estimates is used to construct empirical covariance matrices, which provide finite-sample approximations of the true sampling covariances of the estimators. The empirical counterpart of the theoretical comparison is then obtained by examining the difference between these covariance matrices and by evaluating its spectral properties, in particular through the sign of its smallest eigenvalue. Across all scenarios considered, the empirical results are consistent with the ordering predicted by Proposition 2. In particular, the empirical difference $\widehat{\Sigma}^{R}-\widehat{\Sigma}^{O}$ is found to be positive definite in all tested configurations.
From the standpoint of parameter estimation accuracy, the offset approach may be preferable to the ratio approach in finite samples, as it consistently yields smaller empirical variances for the estimator of $\beta$. This empirical finding is in line with the asymptotic ordering established in Proposition 2.
4.3 Sum of Individual Observed Gaps Comparison
In the context of automobile insurance pricing, one of the primary objectives is not only parameter estimation but also the accurate determination of premiums. From this perspective, we instead focus on proximity to financial balance as a key criterion for evaluating the different approaches. As discussed in Denuit et al. (2024), financial balance broadly refers to the situation in which the sum of observed total losses coincides with the sum of premiums computed under a given pricing method. Since this balance is not exactly achieved under either approach due to the use of a logarithmic link function in premium modeling, financial balance naturally emerges as an important criterion to consider when selecting between the two approaches.
We consider $n$ independent contracts with observed loss costs $y_{i}=t_{i}z_{i}$ for $i=1,\ldots,n$, where $z_{i}$ denotes the loss cost per unit of risk exposure $t_{i}$. We define the individual observed gap as the difference between the observed loss cost and the estimated premium. For contract $i$, these gaps are denoted by $\Delta_{i}^{O}$ and $\Delta_{i}^{R}$ under the offset and ratio approaches, respectively:
$\Delta_{i}^{O}=t_{i}\bigl(z_{i}-\widehat{\zeta}_{i}^{O}\bigr),\qquad\Delta_{i}^{R}=t_{i}\bigl(z_{i}-\widehat{\zeta}_{i}^{R}\bigr),$